OpenStudy (anonymous):

Use l'hopitals to find the limit: lim x->0 tanx/x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sec^2(x) }{ 1 }\]

OpenStudy (anonymous):

I get 0/0 so I need to take the d/dx of the above

OpenStudy (anonymous):

do I have to use the quotient rule ?

OpenStudy (anonymous):

\[=\lim_{x \rightarrow 0}\frac{ 1 }{ \cos^2(x) }\]

OpenStudy (anonymous):

\[\frac{ \frac{ d }{ dx }f(x) }{ \frac{ d }{ dx } g(x)}\]

OpenStudy (anonymous):

YES that's what I was missing, that is the L'hop rule... I was trying to take the derivative as one function. Thank you.

OpenStudy (anonymous):

A medal is appreciated

OpenStudy (anonymous):

Ok so then I get 0 / 1 so i take the derivative again?